OpenStudy (ayonnaleflore):

Which one of the following sets of numbers represents the lengths of the sides of a right triangle? 9, 40, 42 5, 10, 15 13, 35, 37 6, 8, 10

OpenStudy (dayakar):

find squares of all three sides then check sum of any two sides squares equals to third side square

OpenStudy (ayonnaleflore):

im confused

OpenStudy (dayakar):

6,8,10 6^2 = 36 8^2 =64 10^2 = 100 6^2 + 8^2 = 10^2 there fore 6,8,10 are sides of right triangle check other numbers

OpenStudy (dayakar):

is it clear

OpenStudy (ayonnaleflore):

no i still dont understad